OpenStudy (anonymous):

The radius of the base of a cylinder is 38 mm and its height is 51 mm. Find the surface area of the cylinder in terms of pi. A. 6726 pi mm2 B. 6853 pi mm2 C. 6764 pi mm2 D. 6713 pi mm2

OpenStudy (anonymous):

Awnsers C. Area of the two circles, one for the top and one for the bottom. And then the surface Area of the outside Surface. So for the circles, since theres 2 , 2*(pi*r^2) = 2888pi and for the outer surface = the circumfrence of the base * the height. = 2*pi *r * h = 2*pi*38*51 = 3876pi Add the two and you get , 6764 pi

OpenStudy (anonymous):

thank you , would you mind helping me with 2 more problems? :)

OpenStudy (anonymous):

No probs. And Sure but I cant do too many, I gotta study for exams as well. But yea post them up.

OpenStudy (anonymous):

okay thanks :) would you like me to post them in a new one or this one?

OpenStudy (anonymous):

Um just post here I guess, I keep getting notifications for it.

OpenStudy (anonymous):

Find the volume of the square pyramid shown. Round to the nearest tenth if necessary. A. 600 ft3 B. 33.3 ft3 C. 1200 ft3 D. 400 ft3

OpenStudy (anonymous):

So for a Pyramid the volume is the area of the base * height * 1/3. Here the base is a square so the area for the base is simply 10^2 = 100. You'll need to find the hieght using Right Angle Triangles.  So here h is square root of (13^2 - 5^2) using the pythagarous theorem. We get 12. Now the volume should be 100*12*(1/3) = 400

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

yes :) i understand i had a problem similar to this earlier but i forgot to use the pythagorean therom , opps :x , but my last problem is .. Find the surface area of the cylinder in terms of pi. A. 504 pi cm2 B. 333 pi cm2 C. 382.5 pi cm2 D. 211.5 pi cm2

OpenStudy (anonymous):

Allright for this question you just have to think about all the surfaces for the shape. So theres two circles on the top and bottom, If this is a closed cylinder and then theres the outside curve. Now the area for the two circles is 2*(pi*r^2) = 81*pi/2 And for the outside curve, again you just think of it as the circumfrence of the circle being extended. We then get 2*pi*r* the height. = 171*pi. Then add the two up. you get, 211.5 pi cm^2

OpenStudy (anonymous):

omg , thank you soo much! now only if i can remember my formula's , thank you =)

OpenStudy (anonymous):

Yup no problem, you should be fine. good luck!

OpenStudy (anonymous):

same to you on your exams!

OpenStudy (anonymous):

Thanks.